OpenStudy (anonymous):

Write the decimal as a fraction or a mixed number in simplest form. 3.44444444

OpenStudy (campbell_st):

let x = 3.444.... let 10 x = 34.44444. then its 10x - x this eliminates the recurring decimal 10x = 34.4444444 - x = 3.444444 ----------------------- 9x = 31 divide both sides by 9 x = 31/9 or x = 3 4/9 hope this helps